OpenStudy (anonymous):

how do I eliminate the parameter to find a Cartesian equation with the given x=tan(Θ) and y=cot(Θ)?

OpenStudy (anonymous):

we know that tan(t)=1/cot(t) that implies: x=1/y --> xy=1